# Large Language Model-Driven Closed-Loop UAV Operation with Semantic Observations

### v4

@@ Abstract
Index Terms:
UAV, Large Language Model, Robotics, IoT.
@@ I Introduction
Uavs, or drones, are increasingly being integrated into IoT ecosystems for various applications, such as surveillance and monitoring, smart cities, and infrastructure inspection [1]. Their high mobility and rich sensing capabilities make them an ideal candidate for data collection, communication, and service delivery tasks in these applications. Given a UAV task, the traditional pipeline requires a specialized engineer to write the operation code based on the task description and then deploy the code on the device for execution, as shown in Fig. 1(a). Such a process requires substantial domain knowledge and can also be time-consuming. With LLMs’ exceptional semantic understanding and context generation capabilities, recent research has demonstrated the potential of leveraging LLM to facilitate the operation code generation process for a given robotic task [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. As illustrated in Fig. 1(b), these LLM-driven approaches enable the generation of UAV operation code with natural language tasks, which significantly simplifies the code generation and also brings the potential of automated UAV task deployment.
@@
Despite the benefits brought by LLM-driven UAV operation code generation, it also raises concerns about the reliability of LLM-generated code due to the challenges faced by LLM in terms of logical reasoning and complex decision-making [12]. Different from text generation applications, where semantic-level accuracy is sufficient, the generation of UAV operation code requires the precise execution of sequential actions. Erroneous or invalid operation code can lead to unsafe UAV behavior and pose safety and security risks to the public, such as UAV crashes.
@@
To enhance the reliability of LLM in translating natural language tasks into operation code, techniques such as prompt engineering [13] and few-shot learning [14] have been widely adopted when designing LLM-driven operation code generation research for robots, including UAVs, and have demonstrated promising results. However, a set of existing approaches [2, 3, 4, 5] adopt an open-loop design that lacks mechanisms to verify the correctness of the operation code generated by LLMs. Such open-loop designs can leave questions such as cumulative error and temporal action consistency unsolved [15], and hence affect the reliability of their generated code, especially when handling complex tasks. Recently, closed-loop designs have been proposed in LLM-driven robotic operation code generation [6, 7, 8, 9, 10, 11], which enable evaluation of generated code and provide feedback for refinement, thereby steering robot execution towards its task objectives. Nevertheless, these approaches still face limitations from different aspects, including:
@@
  • Using numerical state observations for evaluation: approaches proposed in ref [6, 7] use numerical state observations in scenarios for evaluation and feedback generation. Due to the relatively limited numerical reasoning and processing capabilities of LLMs [16], an LLM-driven evaluation can overlook the latent semantics of robot actions and fail to achieve a comprehensive assessment of its movements and dynamics, leading to unreliable code refinement feedback.
@@
  • Physical execution for evaluation: closed-loop approaches, such as ref [8, 9, 7], employ physical execution of LLM-generated code to generate state observations for evaluation and code refinement. Although these approaches can be acceptable for robots that are not likely to cause physical damage during the refinement, their adoption in the context of UAV operation becomes risky because the physical execution of UAVs with code that potentially contains errors for refinement can easily lead to UAV crashes and safety concerns [17].
@@
  • Self-refinement using a single LLM: existing closed-loop approaches [10, 11] that a single LLM can be used for both code generation, evaluation, and refinement. However, self-evaluation may amplify its errors and lead to degradation in task success rate [18], which is also validated in the experimental evaluation of this work.
@@
To overcome these limitations, this paper proposes a novel closed-loop framework to further enhance the reliability of LLM-driven UAV operation code generation. Specifically, our framework adopts a duo-LLM design, in which one LLM is configured as the Code Generator and the other one is configured as the Evaluator. To compensate for LLM’s limitation in numerical reasoning, we propose a state transformation algorithm that converts numerical state observations into semantic observations by embedding UAV dynamics within the context of semantic trajectory observations during closed-loop evaluation. In addition, our framework enables simulation-based evaluation and refinement of LLM-generated UAV operation code before physical deployment. This feature enables efficient refinement for UAV operation code generation, and more importantly, eliminates the safety risks raised by potential incorrect code execution during refinement with physical UAV execution. Align with existing research [6, 7, 8, 9, 10, 11], this paper also focuses on using LLMs to generate operation code for deployment and task execution with the complete task description provided, but not real-time interactive UAV controls.
@@
The high-level design of our framework is illustrated in Fig. 2. In our framework, the LLM configured as the Code Generator synthesizes or refines UAV operation code based on task description or evaluation feedback, and the LLM configured as the Evaluator assesses code performance and provides feedback to guide code refinement. The process unfolds as follows: 1) The Code Generator first produces an initial code based on the task description; 2) The code is then execute in simulation to produce raw numerical state observation, which are subsequently transformed into a semantic observations; 3) The Evaluator evaluates the semantic observations against the original task specification and returns evaluation feedback; 4) If the evaluation confirms that the observations has met the task objectives, the code will deploy on the UAV for task execution. Otherwise, the evaluation feedback is returned to the Code Generator for iterative refinement.
@@
To evaluate the effectiveness of our proposed framework, we conducted extensive experiments on UAV operation tasks with varying levels of complexity. The experimental results show that our framework achieves a $100\%$ success rate for basic tasks. For advanced tasks with a higher complexity, our framework achieves a $98.5\%$ completeness and an $85\%$ success rate, i.e., $98.5\%$ of required actions in all evaluated tasks are completed correctly and $85\%$ of tasks are entirely completed without any error. Compared with baseline approaches [5, 10, 7], our framework greatly improves both success rate and completeness when the task complexity increases.
@@ II-A LLM for Robotics
Continual advancements in LLMs [19, 20] have demonstrated their exceptional capabilities in semantic understanding and context generation, making them promising when integrating with various robotics tasks. Research works from Microsoft [2] and Google [21] demonstrated how publicly accessible LLMs can be leveraged to generate operation code for various robotic tasks, including UAVs. Since then, a set of research works [3, 4, 5, 6, 7, 8, 9, 10, 11] have been proposed to enhance the quality and reliability of LLM-driven robotic tasks, especially with the integration of prompt engineering techniques [13] that help steer LLMs’ behavior and align their outputs with the tasks’ objectives.
@@
In ref [3], an end-to-end LLM framework is proposed that enables semantic planning with applications in quadcopter drones. This paper also introduced a specific programming language (MiniSpec) to improve the code generation efficiency of LLMs. In ref [4], a prompt-engineered robot control solution is proposed that synthesizes input from both LLMs and humans. This work aims to have the robotic system behavior programmed by emphasizing prompting LLMs and avoiding manipulating code. Recently, a prompt framework with enhanced reasoning, named GSCE [5] was proposed to better support LLM-driven UAV control. GSCE integrates domain-specific guidelines, constraints, and examples into its system prompts when configuring the LLM, and hence improves its reliability of constraint-compliant code generation. While these methods all improve the LLM-driven code generation for robotic tasks from different aspects, they all adopt an open-loop design framework without error correction and adjustment for the generated code. Such a design can lead to potential cumulative errors and temporal action consistency issues [15] in the code generation.
@@
To address the concerns from the open-loop design, different closed-loop refinement strategies are increasingly adopted by recent research towards LLM-driven code generation for robotic tasks. In ref [6, 7], numerical observation of the robot’s execution status, such as actuator reading, sensor reading, and system information (e.g., location), is utilized to construct the feedback for LLMs. These numerical observations-based feedback can positively contribute to the quality of LLM-generated operation code, however, their contribution is restricted when trying to assess the latent semantics of robot actions for a comprehensive assessment. This is because LLMs have relatively limited numerical reasoning and processing capabilities [16]. In addition, to obtain accurate observations and leverage them for the refinement of LLM-generated code, executing the code with physical experiments is employed for feasible robotic tasks. For example, CoPAL [8] employs physical error feedback and run-time error feedback to refine the code when errors occur during physical execution. In ref [9], precondition errors are integrated into a prompt that aims to repair plans when the robot fails skill execution. While evaluating code with physical experiments can provide accurate observations of errors and misalignment between LLM-generated code and task objectives that enable the construction of an effective refinement plan, it does not fit when moving to the context of many UAV tasks. Physical execution of code that may contain errors, especially before any refinement, has a high chance of causing the UAV to crash [17]. This not only increases the hardware cost for the experiment but also can lead to safety concerns. Therefore, our proposed framework enables the execution of LLM-generated UAV operation code in the simulation environment to resolve the safety concerns in the refinement process. Moreover, the state transformation design in our framework enables semantic observations of the execution rather than numerical-only observation, and hence empowers LLMs in the code evaluation and feedback generation for refinement.
@@
In recent research, other strategies to further improve the performance of closed-loop design have also been explored. For example, ref [22] utilizes human feedback to revise the code when misalignment in the generated plan is detected. Although human feedback provides high-quality input for refinement, it limits the automation of the system and the performance depends on the expertise of the user. In addition, existing research [6, 7, 8, 9] separates the configuration of code generator and evaluator on different LLMs. This design aims to minimize the amplification of errors and bias [18] that can be caused by using a single LLM for both code generation and evaluation in previous research [10, 11]. In our framework, we also use two separate LLMs and have customized configurations to improve their capabilities as the code generator and evaluator, respectively.
@@ II-B Prompt Engineering
Recent studies have increasingly employed prompt engineering [13] to improve the reliability of LLM-driven robotic systems. For instance, GSCE proposes a prompt framework as a system prompt to improve the reliability of LLM-driven UAV control code generation [5]. Prompt engineering typically leverages two core techniques: in-context learning [23] and Chain-of-Thought (CoT) reasoning [24]. In-context learning enables the LLM to identify patterns from a limited set of examples and apply them to new situations. This facilitates the generation of control code that adheres to robot policy and how to ground task descriptions from a few examples [21, 25]. On the other hand, CoT prompts the LLM to articulate intermediate inference steps, which is valuable for robotic tasks requiring sequential, stepwise decision-making. By decomposing a task into steps, CoT encourages the production of code that aligns with each stage of the intended action plan. Prior studies have embedded CoT within the examples to guide LLMs through reasoning step-by-step via in-context learning [26, 25]. In this study, we integrate both CoT and in-context learning strategies to enhance the quality of code generation and evaluation.
@@ III Problem Formulation
Given a UAV operation task description, the framework proposed in this paper aims to use LLMs to reason through the task description and generate accurate operation code after observation, evaluation, feedback, and refinement. Specifically, we formulate our problem by extending the LLM-driven UAV operation problem in [5] as $P=\langle s_{0},S,A,T,G\rangle$, where $s_{0}$ is the initial state, $S=(s_{0},s_{1},s_{2},\dots,s_{n})$ is the finite set of discrete states encoding the UAV’s spatial position and orientation, $A$ is the finite action (UAV skills) set, $T$ is the deterministic transition, and $G\subseteq S$ is the goal state. For a current state $S$, an action $a\in A$ is selected from the skill set $A$, leading to a deterministic state transition: $T:S\times A\rightarrow S$ that determines the next state. A candidate solution to problem $P$ is a sequence of actions $l=(a_{1},a_{2},\dots,a_{n})$ that control the UAV from the initial state $s_{0}$ through intermediate states until reaching a final state $s_{n}$. The state transitions of $l$ yield state observations $O$, and evaluation $E=\langle O,P\rangle$ assesses whether $O$ satisfies the problem objectives and outputs $F$ as feedback. If $F$ is “YES”, the sequence of actions $l$ is deemed a valid solution solution that flies the UAV from $s_{0}$ to the goal state $(s_{n}\equiv G$); otherwise, $l$ is iteratively refined based on the feedback $F$.
@@ IV-B Code Generation
The LLM-driven code generation in our framework is achieved by configuring a general-purpose LLM agent (code generator) with UAV operation-related system prompts. We adopt prompt engineering strategies proposed by the GSCE framework [5], which have been demonstrated to be effective for enhancing the reasoning capabilities of LLMs in basic UAV operation code generation. Specifically, the system prompts comprise guidelines, skill APIs, constraints, and examples. The guidelines shape the LLM’s role for code generation, the skill APIs empower LLM to employ appropriate functions, the constraints regulate LLM with robotic policies, and the examples are pairs of query-response that serve as illustrations. These strategies serve as the basic building blocks to support the LLM-driven UAV operation code generation. On top of that, our code generation also utilizes constructive feedback from our evaluation design in section IV-D. The feedback helps the code generator obtain a more comprehensive understanding of the dynamics and interconnections among UAV operations implied by the task description, which enhances the LLM’s reasoning for code generation and code refinement.

### v5

@@ I Introduction
Uavs, or drones, are increasingly being integrated into IoT ecosystems for various applications, such as surveillance and monitoring, smart cities, and infrastructure inspection [uav_iot_applications]. Their high mobility and rich sensing capabilities make them an ideal candidate for data collection, communication, and service delivery tasks in these applications. Given a UAV task, the traditional pipeline requires a specialized engineer to write the operation code based on the task description and then deploy the code on the device for execution, as shown in Fig. 1(a). Such a process requires substantial domain knowledge and can also be time-consuming. With LLMs’ exceptional semantic understanding and context generation capabilities, recent research has demonstrated the potential of leveraging LLM to facilitate the operation code generation process for a given robotic task [ChatGPTRobotics, TypeFly, CLEAR, GSCE, incoro, InteractivePlanning, CoPAL, CAPE, SelfRefine, trust]. As illustrated in Fig. 1(b), these LLM-driven approaches enable the generation of UAV operation code with natural language tasks, which significantly simplifies the code generation and also brings the potential of automated UAV task deployment.
@@
Despite the benefits brought by LLM-driven UAV operation code generation, it also raises concerns about the reliability of LLM-generated code due to the challenges faced by LLM in terms of logical reasoning and complex decision-making [UAVsmeetllmsoverviews]. Different from text generation applications, where semantic-level accuracy is sufficient, the generation of UAV operation code requires the precise execution of sequential actions. Erroneous or invalid operation code can lead to unsafe UAV behavior and pose safety and security risks to the public, such as UAV crashes.
@@
To enhance the reliability of LLM in translating natural language tasks into operation code, techniques such as prompt engineering [PromptEngineering] and few-shot learning [few-shot] have been widely adopted when designing LLM-driven operation code generation research for robots, including UAVs, and have demonstrated promising results. However, a set of existing approaches [ChatGPTRobotics, TypeFly, CLEAR, GSCE] adopt an open-loop design that lacks mechanisms to verify the correctness of the operation code generated by LLMs. Such open-loop designs can leave questions such as cumulative error and temporal action consistency unsolved [OpenLoopQuestions], and hence affect the reliability of their generated code, especially when handling complex tasks. Recently, closed-loop designs have been proposed in LLM-driven robotic operation code generation [incoro, InteractivePlanning, CoPAL, CAPE, SelfRefine, trust], which enable evaluation of generated code and provide feedback for refinement, thereby steering robot execution towards its task objectives. Nevertheless, these approaches still face limitations from different aspects, including:
@@
  • Using numerical state observations for evaluation: approaches proposed in ref [incoro, InteractivePlanning] use numerical state observations in scenarios for evaluation and feedback generation. Due to the relatively limited numerical reasoning and processing capabilities of LLMs [NumberUnderstanding], an LLM-driven evaluation can overlook the latent semantics of robot actions and fail to achieve a comprehensive assessment of its movements and dynamics, leading to unreliable code refinement feedback.
@@
  • Physical execution for evaluation: closed-loop approaches, such as ref [CoPAL, CAPE, InteractivePlanning], employ physical execution of LLM-generated code to generate state observations for evaluation and code refinement. Although these approaches can be acceptable for robots that are not likely to cause physical damage during the refinement, their adoption in the context of UAV operation becomes risky because the physical execution of UAVs with code that potentially contains errors for refinement can easily lead to UAV crashes and safety concerns [uav-crash].
@@
  • Self-refinement using a single LLM: existing closed-loop approaches [SelfRefine, trust] that a single LLM can be used for both code generation, evaluation, and refinement. However, self-evaluation may amplify its errors and lead to degradation in task success rate [LLM-bias], which is also validated in the experimental evaluation of this work.
@@
To overcome these limitations, this paper proposes a novel closed-loop framework to further enhance the reliability of LLM-driven UAV operation code generation. Specifically, our framework adopts a duo-LLM design, in which one LLM is configured as the Code Generator and the other one is configured as the Evaluator. To compensate for LLM’s limitation in numerical reasoning, we propose a state transformation algorithm that converts numerical state observations into semantic observations by embedding UAV dynamics within the context of semantic trajectory observations during closed-loop evaluation. In addition, our framework enables simulation-based evaluation and refinement of LLM-generated UAV operation code before physical deployment. This feature enables efficient refinement for UAV operation code generation, and more importantly, eliminates the safety risks raised by potential incorrect code execution during refinement with physical UAV execution. Align with existing research [incoro, InteractivePlanning, CoPAL, CAPE, SelfRefine, trust], this paper also focuses on using LLMs to generate operation code for deployment and task execution with the complete task description provided, but not real-time interactive UAV controls.
@@
The high-level design of our framework is illustrated in Fig. 2. In our framework, the LLM configured as the Code Generator synthesizes or refines UAV operation code based on task description or evaluation feedback, and the LLM configured as the Evaluator assesses code performance and provides feedback to guide code refinement. The process unfolds as follows: 1) The Code Generator first produces an initial code based on the task description; 2) The code is then executed in simulation to produce raw numerical state observation, which are subsequently transformed into semantic observations; 3) The Evaluator evaluates the semantic observations against the original task specification and returns evaluation feedback; 4) If the evaluation confirms that the observations have met the task objectives, the code will deploy on the UAV for task execution. Otherwise, the evaluation feedback is returned to the Code Generator for iterative refinement.
@@
To evaluate the effectiveness of our proposed framework, we conducted extensive experiments on UAV operation tasks with varying levels of complexity. The experimental results show that our framework achieves a $100\%$ success rate for basic tasks. For advanced tasks with a higher complexity, our framework achieves a $98.5\%$ completeness and an $85\%$ success rate, i.e., $98.5\%$ of required actions in all evaluated tasks are completed correctly and $85\%$ of tasks are entirely completed without any error. Compared with baseline approaches [GSCE, SelfRefine, InteractivePlanning], our framework greatly improves both success rate and completeness when the task complexity increases.
@@ II-A LLM for Robotics
Continual advancements in LLMs [deepseek, o3mini] have demonstrated their exceptional capabilities in semantic understanding and context generation, making them promising when integrating with various robotics tasks. Research works from Microsoft [ChatGPTRobotics] and Google [CodeasPolicies] demonstrated how publicly accessible LLMs can be leveraged to generate operation code for various robotic tasks, including UAVs. Since then, a set of research works [TypeFly, CLEAR, GSCE, incoro, InteractivePlanning, CoPAL, CAPE, SelfRefine, trust] have been proposed to enhance the quality and reliability of LLM-driven robotic tasks, especially with the integration of prompt engineering techniques [PromptEngineering] that help steer LLMs’ behavior and align their outputs with the tasks’ objectives.
@@
In ref [TypeFly], an end-to-end LLM framework is proposed that enables semantic planning with applications in quadcopter drones. This paper also introduced a specific programming language (MiniSpec) to improve the code generation efficiency of LLMs. In ref [CLEAR], a prompt-engineered robot control solution is proposed that synthesizes input from both LLMs and humans. This work aims to have the robotic system behavior programmed by emphasizing prompting LLMs and avoiding manipulating code. Recently, a prompt framework with enhanced reasoning, named GSCE [GSCE], was proposed to better support LLM-driven UAV control. GSCE integrates domain-specific guidelines, constraints, and examples into its system prompts when configuring the LLM, and hence improves its reliability of constraint-compliant code generation. While these methods all improve the LLM-driven code generation for robotic tasks from different aspects, they all adopt an open-loop design framework without error correction and adjustment for the generated code. Such a design can lead to potential cumulative errors and temporal action consistency issues [OpenLoopQuestions] in the code generation.
@@
To address the concerns from the open-loop design, different closed-loop refinement strategies are increasingly adopted by recent research towards LLM-driven code generation for robotic tasks. In ref [incoro, InteractivePlanning], numerical observation of the robot’s execution status, such as actuator reading, sensor reading, and system information (e.g., location), is utilized to construct the feedback for LLMs. These numerical observations-based feedback can positively contribute to the quality of LLM-generated operation code, however, their contribution is restricted when trying to assess the latent semantics of robot actions for a comprehensive assessment. This is because LLMs have relatively limited numerical reasoning and processing capabilities [NumberUnderstanding]. In addition, to obtain accurate observations and leverage them for the refinement of LLM-generated code, executing the code with physical experiments is employed for feasible robotic tasks. For example, CoPAL [CoPAL] employs physical error feedback and run-time error feedback to refine the code when errors occur during physical execution. In ref [CAPE], precondition errors are integrated into a prompt that aims to repair plans when the robot fails skill execution. While evaluating code with physical experiments can provide accurate observations of errors and misalignment between LLM-generated code and task objectives that enable the construction of an effective refinement plan, it does not fit when moving to the context of many UAV tasks. Physical execution of code that may contain errors, especially before any refinement, has a high chance of causing the UAV to crash [uav-crash]. This not only increases the hardware cost for the experiment but also can lead to safety concerns. Therefore, our proposed framework enables the execution of LLM-generated UAV operation code in the simulation environment to resolve the safety concerns in the refinement process. Moreover, the state transformation design in our framework enables semantic observations of the execution rather than numerical-only observation, and hence empowers LLMs in the code evaluation and feedback generation for refinement.
@@
In recent research, other strategies to further improve the performance of closed-loop design have also been explored. For example, ref [preemptive] utilizes human feedback to revise the code when misalignment in the generated plan is detected. Although human feedback provides high-quality input for refinement, it limits the automation of the system and the performance depends on the expertise of the user. In addition, existing research [incoro, InteractivePlanning, CoPAL, CAPE] separates the configuration of the code generator and evaluator on different LLMs. This design aims to minimize the amplification of errors and bias [LLM-bias] that can be caused by using a single LLM for both code generation and evaluation in previous research [SelfRefine, trust]. In our framework, we also use two separate LLMs and have customized configurations to improve their capabilities as the code generator and evaluator, respectively.
@@ II-B Prompt Engineering
Recent studies have increasingly employed prompt engineering [PromptEngineering] to improve the reliability of LLM-driven robotic systems. For instance, GSCE proposes a prompt framework as a system prompt to improve the reliability of LLM-driven UAV control code generation [GSCE]. Prompt engineering typically leverages two core techniques: in-context learning [URIAL] and Chain-of-Thought (CoT) reasoning [CoT]. In-context learning enables the LLM to identify patterns from a limited set of examples and apply them to new situations. This facilitates the generation of control code that adheres to robot policy and how to ground task descriptions from a few examples [CodeasPolicies, PromptBook]. On the other hand, CoT prompts the LLM to articulate intermediate inference steps, which is valuable for robotic tasks requiring sequential, stepwise decision-making. By decomposing a task into steps, CoT encourages the production of code that aligns with each stage of the intended action plan. Prior studies have embedded CoT within the examples to guide LLMs through reasoning step-by-step via in-context learning [ISR-LLM, PromptBook]. In this study, we integrate both CoT and in-context learning strategies to enhance the quality of code generation and evaluation.
@@ III Problem Formulation
Given a UAV operation task description, the framework proposed in this paper aims to use LLMs to reason through the task description and generate accurate operation code after observation, evaluation, feedback, and refinement. Specifically, we formulate our problem by extending the LLM-driven UAV operation problem in [GSCE] as $P=\langle s_{0},S,A,T,G\rangle$, where $s_{0}$ is the initial state, $S=(s_{0},s_{1},s_{2},\dots,s_{n})$ is the finite set of discrete states encoding the UAV’s spatial position and orientation, $A$ is the finite action (UAV skills) set, $T$ is the deterministic transition, and $G\subseteq S$ is the goal state. For a current state $S$, an action $a\in A$ is selected from the skill set $A$, leading to a deterministic state transition: $T:S\times A\rightarrow S$ that determines the next state. A candidate solution to problem $P$ is a sequence of actions $l=(a_{1},a_{2},\dots,a_{n})$ that control the UAV from the initial state $s_{0}$ through intermediate states until reaching a final state $s_{n}$. The state transitions of $l$ yield state observations $O$, and evaluation $E=\langle O,P\rangle$ assesses whether $O$ satisfies the problem objectives and outputs $F$ as feedback. If $F$ is “YES”, the sequence of actions $l$ is deemed a valid solution that flies the UAV from $s_{0}$ to the goal state $(s_{n}\equiv G$); otherwise, $l$ is iteratively refined based on the feedback $F$.
@@ IV-B Code Generation
The LLM-driven code generation in our framework is achieved by configuring a general-purpose LLM agent (code generator) with UAV operation-related system prompts. We adopt prompt engineering strategies proposed by the GSCE framework [GSCE], which have been demonstrated to be effective for enhancing the reasoning capabilities of LLMs in basic UAV operation code generation. Specifically, the system prompts comprise guidelines, skill APIs, constraints, and examples. The guidelines shape the LLM’s role for code generation, the skill APIs empower LLM to employ appropriate functions, the constraints regulate LLM with robotic policies, and the examples are pairs of query-response that serve as illustrations. These strategies serve as the basic building blocks to support the LLM-driven UAV operation code generation. On top of that, our code generation also utilizes constructive feedback from our evaluation design in section IV-D. The feedback helps the code generator obtain a more comprehensive understanding of the dynamics and interconnections among UAV operations implied by the task description, which enhances the LLM’s reasoning for code generation and code refinement.